If the time period of simple pendulum is 2 s. Calculate time that its Bob takes to move from extreme position to mean position.
1
Expert's answer
2018-04-30T07:56:07-0400
The time needed that Bob takes to move from maximum position to mean position
∆t=T/4
So
∆t=2/4=0.5 s